ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 200, 2014

Abels, H. and A. Manoussos Linear semigroups with coarsely dense orbits

327

Alesker, S. Corrigendum to “A Fourier type transform on translation invariant valuations on convex sets”

507

Archbold, R. J. and D. W. B. Somerset Separation properties in the primitive ideal space of a multiplier algebra

389

Argyros, S. A., A. Manoussakis and A. Pelczar-Barwacz Quasi-minimality and tightness by range in spaces with unconditional basis

19

Barbosa, E., R. Pina and K. Tenenblat On gradient Ricci solitons conformal to a pseudo-Euclidean space

213

Blagojević, P. V. M. and G. M. Ziegler Convex equipartitions via Equivariant Obstruction Theory

49

Bodirsky, M. and M. Pinsker Minimal functions on the random graph

251

Cantón, A. and J. Qu A note on asymptotic values of quasiregular maps

343

Chuaqui, M., H. Hamada, R. Hernández and G. Kohr Pluriharmonic mappings and linearly connected domains in ℂ^{n}

489

Clop, A. and D. A. Herron Mappings with finite distortion in L_{loc}^{p}: modulus of continuity and compression of Hausdorff measure

225

Coskun, I. Rigidity of Schubert classes in orthogonal Grassmannians  85
Gerstenhaber, M. Self-dual and quasi self-dual algebras  193

Hamada, H. See M. Chuaqui

Hamilton, M. J. D. The minimal genus problem for elliptic surfaces  127

Hernández, R. See M. Chuaqui Herron, D. A. See A. Clop Hillman, J., D. Kochloukova and I. Lima Pro-p completions of

Poincaré duality groups

1

Kloeckner, B. R. Yet another short proof of Bourgain’s distortion estimate for embedding of trees into uniformly convex Banach spaces

419

Kochloukova, D. See J. Hillman Kohr, G. See M. Chuaqui Lascar, R. and J. Nourrigat Gårding inequality in large dimen-

sion

79

Lima, I. See J. Hillman Manoussakis, A. See S. A. Argyros Manoussos, A. See H. Abels Marshall, S. Triple product L-functions and quantum chaos on

SL(2, ℂ) 423

Nourrigat, J. See R. Lascar Pelczar-Barwacz, A. See S. A. Argyros Pina, R. See E. Barbosa Pinsker, M. See M. Bodirsky Qu, J. See A. Cantón Regev, A. Kronecker multiplicities in the (k, ℓ) hook are polynomially

bounded

39

Schmidt, R. On the subgroup lattice index problem in finite groups  141
Simis, A. and S. O. Tohǎneanu Homology of homogeneous divisors  449

Somerset, D. W. B. See R. J. Archbold Srivastava, R. K. Real analytic expansion of spectral projections and

extension of the Hecke–Bochner identity

171

Steele, A. Nonassociative cyclic algebras  361

Tenenblat, K. See E. Barbosa Tohǎneanu, S. O. See A. Simis

Zernik, A. Open conformal cobordisms  297

Ziegler, G. M. See P. V. M. Blagojević